I am not to sure of myself in Algebra and need to graph something...I have tried but as I said I am not sure can you please tell me how to graph the inequality
x + y = 3

Answer by mananth(16946)

x + y = 3
this is an equation of a line
you need to generate values of y from various values of x
let x = 0,
x+y =3
0+y =3 so y=3
One point (0,3)
if x = 2 then 2+y =3 OR y = 3-2 =1
So the point is (2,1)
when y = 0, x=3. third point (3,0)
Now you have 3 points . plot them on the graph and draw the line graph.

m.ananth@hotmail.ca